OpenStudy (anonymous):

Write an equation of a line perpendicular to line CD below in slope-intercept form that passes through the point (−1, 6). Line CD is shown. C is at 1, 1. D is at 3, 5 y = −0.5x − 5.5 y = −0.5x + 5.5 y = 2x + 13 y = 2x − 13

OpenStudy (anonymous):

First find the gradient of line CD

OpenStudy (anonymous):

But what's the formula

OpenStudy (nurali):

You have a point which is: (−1,6) Remember that points are written as: (x,y) Now it is the plug and play game: y=−0.5x+5.5 x=−1,y=6 6=−0.5(−1)+5.5 6=0.5+5.5 6=6 6=6 is true, thus your answer is: A) y=−0.5x+5.5

OpenStudy (anonymous):

Okay thank you @Nurali

OpenStudy (nurali):

My Pleasure.